OpenStudy (anonymous):

Given a triangle, with sides 3, 4 and 5, determine if it is a right triangle. A. What is the formula you will use to solve this problem? B. Show all your steps you used to determine if this problem is a right triangle?

OpenStudy (anonymous):

well try drawing it and finding angles.

OpenStudy (anonymous):

also welcome to openstudy

OpenStudy (abb0t):

HINT: pythagorean

OpenStudy (anonymous):

hmm i can try the pathagorian therom.

OpenStudy (anonymous):

and thanks ronncc

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

also just for spelling's sake *pythagorean - from pythagoras ~ a greek polymath

OpenStudy (anonymous):

thanks for auto corret to X]

OpenStudy (anonymous):

..................... *correct *too

OpenStudy (jiteshmeghwal9):

(p)^2+{b}^2=(h)^2 & remember put the longer side instead of h since h stands for hypotenuse.

OpenStudy (jiteshmeghwal9):

(3)^2+(4)^2=(5)^2 now it will be easy to you

OpenStudy (anonymous):

ok thanks everyone

OpenStudy (jiteshmeghwal9):

yw :)